# competing interactions; square well; square shoulder; discontinuous structural crossover transitions; Fisher–Widom line; Rational-Function Approximation

### v1

@@ 2.1 Theoretical background
2.2 $\epsilon_{1}=\epsilon_{2}>0$. Influence of $\lambda_{2}$ on the DOC line
In the case $\epsilon_{1}=\epsilon_{2}>0$, the potential in Equation (1) simplifies to a hard core plus a square shoulder of width $\lambda_{2}-1$. The key question we aim to address is how the DOC line is affected as $\lambda_{2}$ is varied.
As shown in Figure 1, the DOC line exhibits an intricate behavior as $\lambda_{2}$ varies. For $\lambda_{2}=2$ and $\lambda_{2}=1.9$, distinct DOC lines emerge, each starting at $\rho^{*}=\lambda_{2}^{-1}$ in the low-temperature region and shifting toward lower densities as the temperature increases. When $\lambda_{2}=1.8$, the DOC line intersects with a DOC loop at $\rho^{*}\approx 0.18$ and $T^{*}\approx 35$. Inside the loop, the oscillation frequency reaches values $\omega/2\pi\approx 3/\lambda_{2}$, significantly larger than outside the loop. The intersection between the DOC line and the DOC loop acts as a triple point, where three distinct complex poles share the same real part, $\zeta$. An additional DOC arc appears, extending between $\rho^{*}=\lambda_{2}^{-1}$ and $\rho^{*}=1$ in the low-temperature region, within which $\omega$ reaches even higher values ($\omega/2\pi\approx 5=9/\lambda_{2}$, see Appendix A.3.2) than inside the DOC loop. For $\lambda_{2}=1.7$, the loop expands, shifting toward higher densities and lower temperatures, while the DOC arc broadens. Below $T^{*}=0.01$ (not shown in the figure), an inner arc emerges, which is absent in the case $\lambda_{2}=1.8$. In the region between the inner and outer arcs for $\lambda_{2}=1.7$, $\omega/2\pi\approx 5/\lambda_{2}$, whereas $\omega/2\pi\approx 10=17/\lambda_{2}$ within the inner arc. As $\lambda_{2}$ further decreases to $1.6$, the original DOC line vanishes, with the loop and outer arc merging into a more complex DOC region (where $\omega/2\pi\approx 3/\lambda_{2}$) and the inner arc region (where $\omega/2\pi\approx 5=8/\lambda_{2}$) growing. At $\lambda_{2}=1.5$ only the inner arc persists, with $\omega/2\pi\approx 2=3\lambda_{2}$ within. This evolution illustrates an increasingly complex pattern of structural transitions as the DOC line transforms with decreasing $\lambda_{2}$.
@@
Figure 1: DOC lines on the $T^{*}$ vs $\rho^{*}$ plane for the 1D case with $\epsilon_{1}=\epsilon_{2}$: (a) $\lambda_{2}=2,1.9$, (b) $\lambda_{2}=1.8,1.7$, and (c) $\lambda_{2}=1.6,1.5$. Insets display the angular frequency of the asymptotic oscillations of $h(r)$ as a function of $T^{*}$ for (a) $\lambda_{2}=1.9$ at $\rho^{*}=0.5$, (b) $\lambda_{2}=1.7$ at $\rho^{*}=0.5$, and (c) $\lambda_{2}=1.6$ at $\rho^{*}=0.8$.
The insets in Figure 1 illustrate the temperature dependence of $\omega$ at several densities and values of $\lambda_{2}$. In the insets of Figures 1(a) and 1(b), $\omega$ transitions from $\lambda_{2}^{-1}\omega_{\text{HR}}(\rho^{*}\lambda_{2})$ at low $T^{*}$ to $\omega_{\text{HR}}(\rho^{*})$ at high $T^{*}$. In the inset of Figure 1(a), a single discontinuous shift is observed as the DOC line is traversed. However, in the inset of Figure 1(b), two distinct discontinuous jumps in $\omega$ occur as the DOC loop is crossed. The inset of Figure 1(c) shows two discontinuous drops in $\omega$ when crossing the DOC inner and outer arcs.
@@
Figure 2: Logarithmic plot of $|h(r)|$ for large $r$ in the 1D case $\epsilon_{1}=\epsilon_{2}$, $\lambda_{2}=1.7$, for the following states: (a) $(\rho^{*},T^{*})=(0.55,1)$, (b) $(\rho^{*},T^{*})=(0.55,3)$, (c) $(\rho^{*},T^{*})=(0.55,5)$, and (d) $(\rho^{*},T^{*})=(0.8,0.05)$.
As further confirmation of the results presented in Figure 1, we numerically inverted the Laplace transform given by Equation (3) using the method described in Yuste (2023) to obtain $h(r)$. The results for $\lambda_{2}=1.7$ and four representative states are shown in Figure 2. In Figures 2(a)–(c), we fix the density $\rho^{*}$ and examine a temperature ($T^{*}=1$) below the loop, a temperature ($T^{*}=3$) inside the loop, and a temperature ($T^{*}=5$) above the loop. The corresponding leading poles are $(\zeta,\omega)=(1.254,3.400)$, $(\zeta,\omega)=(1.881,10.564)$, and $(\zeta,\omega)=(1.696,4.687)$, respectively, which align fully with the damped oscillatory behavior observed in Figures 2(a)–(c). As a representative state within the region between the inner and outer arcs, we considered $\rho^{*}=0.8$ and $T^{*}=0.05$ in Figure 2(d), for which $(\zeta,\omega)=(0.131,18.233)$.
2.3 $\lambda_{1}=1.35$, $\lambda_{2}=1.7$. Influence of $\epsilon_{1}/\epsilon_{2}$ on the DOC line
If $\epsilon_{1}\neq\epsilon_{2}$, both $\lambda_{1}$ and $\lambda_{2}$ become relevant parameters. For symmetry reasons, we will choose $\lambda_{1}-1=\lambda_{2}-\lambda_{1}$, so that both sections have the same width. As mentioned in Section 1, and to maintain concreteness, we will henceforth set $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$.
@@
Figure 3: DOC lines on the $T^{*}$ vs $\rho^{*}$ plane for the 1D case with $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$: (a) $\epsilon_{1}/\epsilon_{2}=1,0.5$, (b) $\epsilon_{1}=0$, and (c) $\epsilon_{1}/\epsilon_{2}=-0.5,-1$. Insets display the angular frequency of the asymptotic oscillations of $h(r)$ as a function of $T^{*}$ for (a) $\epsilon_{1}/\epsilon_{2}=0.5$ at $\rho^{*}=0.45$, (b) $\epsilon_{1}=0$ at $\rho^{*}=0.22$, and (c) $\epsilon_{1}/\epsilon_{2}=-1$ at $\rho^{*}=0.22$.
Figure 3(a) shows the DOC line, loop, and arc for $\epsilon_{1}/\epsilon_{2}=1$ [also displayed in Figure 1(b)] and for $\epsilon_{1}/\epsilon_{2}=0.5$. In the latter case, the loop expands and shifts up and to the left, while the arc moves downward. For a fluid with $\epsilon_{1}=0$ [Figure 3(b)], the DOC line appears at a density below $\lambda_{2}^{-1}$, with the loop evolving into a lobe that emerges from the vertical axis at $\rho^{*}=0$. In Figure 3(c), a short DOC line forms at zero density when $\epsilon_{1}/\epsilon_{2}=-0.5$, but it vanishes when $\epsilon_{1}/\epsilon_{2}=-1$. As $\epsilon_{1}/\epsilon_{2}$ becomes increasingly negative, the DOC lobe progressively contracts, moving up and to the left until it eventually disappears (not shown). The insets in Figure 3 show the oscillation frequency $\omega$ as a function of $T^{*}$ at selected values of density and the energy ratio $\epsilon_{1}/\epsilon_{2}$.
2.4 $\lambda_{1}=1.35$, $\lambda_{2}=1.7$, $\epsilon_{1}<0$. Influence of $\epsilon_{1}/\epsilon_{2}$ on the FW line
@@
We now consider the case $\epsilon_{1}<0$, where a genuine competition arises between the attractive square well with depth $|\epsilon_{1}|$ and the repulsive barrier of height $\epsilon_{2}$. As demonstrated in Appendix A.1, real poles of $G(s)$ may exist. If one of these real poles becomes dominant, the asymptotic decay of $h(r)$ is monotonic, and, as mentioned earlier, an FW line emerges, marking the abrupt transition between monotonic and oscillatory decay. However, a DOC line may still occur, as exemplified by Figure 3(c).
Figure 4: (a) FW lines on the $T^{*}$ vs $\rho^{*}$ plane for the 1D case $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$ with, from bottom to top, $\epsilon_{1}/\epsilon_{2}=-0.5$, $-1$, $-2$, $-4$, $-10$, and $-100$. (b) Same as in panel (a), except that now the vertical axis represents the scaled temperature $T_{1}^{*}=k_{B}T/|\epsilon_{1}|=T^{*}\epsilon_{2}/|\epsilon_{1}|$. The dotted curve is the FW line for a pure square-well fluid ($\epsilon_{2}=0$). Note that in panel (b) the curves corresponding to $\epsilon_{1}/\epsilon_{2}=-0.5$ and $-1$ are indistinguishable.
@@
The results for various values of $\epsilon_{1}/\epsilon_{2}<0$ are presented in Figure 4(a). A comparison of the DOC lines in Figure 3(c) for $\epsilon_{1}/\epsilon_{2}=-0.5$ and $-1$ with the corresponding FW lines in Figure 4(a) shows that the FW lines emerge at significantly lower values of $T^{*}$ and span a broader range of densities.
@@
The strong sensitivity of the FW lines to the values of $\epsilon_{1}/\epsilon_{2}$, as seen in Figure 4(a), is significantly reduced when temperature is scaled by the well depth $|\epsilon_{1}|$, i.e., $T_{1}^{*}=k_{B}T/|\epsilon_{1}|=T^{*}\epsilon_{2}/|\epsilon_{1}|$. This rescaling is applied in Figure 4(b), which also includes the FW line for a pure square-well fluid ($\epsilon_{2}/|\epsilon_{1}|\to 0$). For the pure square-well fluid, the FW line approaches $T_{1}^{*}\to\infty$ as $\rho^{*}\to 0$ (following a power law). However, introducing a repulsive barrier of height $\epsilon_{2}$ causes the FW line to bend at low densities, even when $\epsilon_{2}/|\epsilon_{1}|=10^{-2}$.
3 The 3D system. RFA results
@@ 3.1 Theoretical background
In this section, we provide a brief account of the main outcome of the RFA approach when the intermolecular potential in 3D is of the form of Equation (1). The detailed derivation may be found in References Santos et al. (2012, 2013); Yuste et al. (2022). We begin by considering a function $G(s)$, which is distinct from its 1D counterpart. This function represents the Laplace transform of $rg(r)$, specifically
@@
We next define an auxiliary function $\Phi(s)$ directly related to $G(s)$ through
@@
Note that Equations (12)–(14) are formally exact. Finally, to construct our RFA, we assume the following rational-function approximate form for $R_{j}(s)$:
@@
The approximation (15) contains nine parameters to be determined by application of certain constraints Santos et al. (2012). The expressions for those nine coefficients are presented in Appendix B.
Once again, the total correlation function $h(r)$ can be expressed in terms of the nonzero poles of $G(s)$, which, in principle, form an infinite set. These poles may be either real or occur in complex-conjugate pairs. Their locations depend on the thermodynamic state, and as before, the pole with the real part closest to zero dictates the asymptotic behavior of the total correlation function for a given state. The 3D analogs of Equations (5) and (6) are, respectively,
@@
In the context of the RFA, Equations (13)–(15) imply that the complex poles satisfy the following set of coupled equations:
@@
3.2 $\epsilon_{1}=\epsilon_{2}>0$. Influence of $\lambda_{2}$ on the DOC line

### v2

@@ 2.1 Theoretical Background
2.2 $\epsilon_{1}=\epsilon_{2}>0$: Influence of $\lambda_{2}$ on DOC Line
In the case $\epsilon_{1}=\epsilon_{2}>0$, the potential in Equation (1) simplifies to a hard core plus a square shoulder of the width $\lambda_{2}-1$.
As shown in Figure 1, the DOC line exhibits an intricate behavior as $\lambda_{2}$ varies. For $\lambda_{2}=2$ and $\lambda_{2}=1.9$, distinct DOC lines emerge, each starting at $\rho^{*}=\lambda_{2}^{-1}$ for the low-temperature region and shifting toward lower densities as the temperature increases. When $\lambda_{2}=1.8$, the DOC line intersects with a DOC loop at $\rho^{*}\approx 0.18$ and $T^{*}\approx 35$. Inside the loop, the oscillation frequency reaches the values $\omega/2\pi\approx 3/\lambda_{2}$, significantly larger than outside the loop. The intersection between the DOC line and the DOC loop acts as a triple point, where three distinct complex poles share the same real part, $\zeta$. An additional DOC arc appears, extending between $\rho^{*}=\lambda_{2}^{-1}$ and $\rho^{*}=1$ for the low-temperature region, within which $\omega$ reaches even higher values ($\omega/2\pi\approx 5=9/\lambda_{2}$; see Appendix A.3.2) than inside the DOC loop. For $\lambda_{2}=1.7$, the loop expands, shifting toward higher densities and lower temperatures, while the DOC arc broadens. Below $T^{*}=0.01$ (not shown in the figure), an inner arc emerges, which is absent in the case $\lambda_{2}=1.8$. In the region between the inner and outer arcs for $\lambda_{2}=1.7$, $\omega/2\pi\approx 5/\lambda_{2}$, whereas $\omega/2\pi\approx 10=17/\lambda_{2}$ within the inner arc. As $\lambda_{2}$ further decreases to $1.6$, the original DOC line vanishes, with the loop and outer arc merging into a more complex DOC region (where $\omega/2\pi\approx 3/\lambda_{2}$) and the inner arc region (where $\omega/2\pi\approx 5=8/\lambda_{2}$) growing. At $\lambda_{2}=1.5$, only the inner arc persists, with $\omega/2\pi\approx 2=3\lambda_{2}$ within. This evolution illustrates an increasingly complex pattern of structural transitions as the DOC line transforms with decreasing $\lambda_{2}$.
@@ 2.2 $\epsilon_{1}=\epsilon_{2}>0$: Influence of $\lambda_{2}$ on DOC Line
Figure 1: DOC lines on the $T^{*}$ vs. $\rho^{*}$ plane for the 1D case with $\epsilon_{1}=\epsilon_{2}$: (a) $\lambda_{2}=2,1.9$, (b) $\lambda_{2}=1.8,1.7$, and (c) $\lambda_{2}=1.6,1.5$. Insets display the angular frequency of the asymptotic oscillations of $h(r)$ as a function of $T^{*}$ for (a) $\lambda_{2}=1.9$ at $\rho^{*}=0.5$, (b) $\lambda_{2}=1.7$ at $\rho^{*}=0.5$, and (c) $\lambda_{2}=1.6$ at $\rho^{*}=0.8$. The circles in panel (b) represent the four states examined in Figure 2 for $\lambda_{2}=1.7$.
The insets in Figure 1 illustrate the temperature dependence of $\omega$ at several densities and values of $\lambda_{2}$. In the insets of Figure 1a,b, $\omega$ transitions from $\lambda_{2}^{-1}\omega_{\text{HR}}(\rho^{*}\lambda_{2})$ at a low $T^{*}$ to $\omega_{\text{HR}}(\rho^{*})$ at a high $T^{*}$. In the inset of Figure 1a, a single discontinuous shift is observed as the DOC line is traversed. However, in the inset of Figure 1b, two distinct discontinuous jumps in $\omega$ occur as the DOC loop is crossed. The inset of Figure 1c shows two discontinuous drops in $\omega$ when crossing the DOC’s inner and outer arcs.
As further confirmation of the results presented in Figure 1, we numerically invert the Laplace transform given by Equation (3) using the method described in Yuste (2023) to obtain $h(r)$. The results for $\lambda_{2}=1.7$ and four representative states are shown in Figure 2. In Figure 2a–c, we fix the density $\rho^{*}$ and examine a temperature ($T^{*}=1$) below the loop, a temperature ($T^{*}=3$) inside the loop, and a temperature ($T^{*}=5$) above the loop. These three states are labeled A–C in Figure 1b, respectively. The corresponding leading poles are $(\zeta,\omega)=(1.254,3.400)$, $(\zeta,\omega)=(1.881,10.564)$, and $(\zeta,\omega)=(1.696,4.687)$, respectively, which align fully with the damped oscillatory behavior observed in Figure 2a–c. As a representative state located between the inner and outer arcs, we select $\rho^{*}=0.8$ and $T^{*}=0.05$ [see label D in Figure 1b]. The corresponding values of the decay parameters are $(\zeta,\omega)=(0.131,18.233)$, as shown in Figure 2d.
@@
Figure 2: A logarithmic plot of $|h(r)|$ for large $r$ in the 1D case $\epsilon_{1}=\epsilon_{2}$, $\lambda_{2}=1.7$, for the following states: (a) $(\rho^{*},T^{*})=(0.55,1)$, (b) $(\rho^{*},T^{*})=(0.55,3)$, (c) $(\rho^{*},T^{*})=(0.55,5)$, and (d) $(\rho^{*},T^{*})=(0.8,0.05)$. These states are labeled A–D in Figure 1b, respectively.
2.3 $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$: Influence of $\epsilon_{1}/\epsilon_{2}$ on DOC Line
If $\epsilon_{1}\neq\epsilon_{2}$, both $\lambda_{1}$ and $\lambda_{2}$ become relevant parameters. For symmetry reasons, we choose $\lambda_{1}-1=\lambda_{2}-\lambda_{1}$ so that both sections have the same width. As mentioned in Section 1 and to maintain concreteness, we henceforth set $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$.
Figure 3a shows the DOC line, loop, and arc for $\epsilon_{1}/\epsilon_{2}=1$ [also displayed in Figure 1b] and for $\epsilon_{1}/\epsilon_{2}=0.5$. In the latter case, the loop expands and shifts up and to the left, while the arc moves downward. For a fluid with $\epsilon_{1}=0$ [Figure 3b], the DOC line appears at a density below $\lambda_{2}^{-1}$, with the loop evolving into a lobe that emerges from the vertical axis at $\rho^{*}=0$. In Figure 3c, a short DOC line forms at very small densities when $\epsilon_{1}/\epsilon_{2}=-0.5$, but it vanishes when $\epsilon_{1}/\epsilon_{2}=-1$. As $\epsilon_{1}/\epsilon_{2}$ becomes increasingly negative, we have observed that the DOC lobe progressively contracts, moving up and to the left until it eventually disappears. The insets in Figure 3 show the oscillation frequency $\omega$ as a function of $T^{*}$ at selected values of density and the energy ratio $\epsilon_{1}/\epsilon_{2}$.
@@
Figure 3: DOC lines on the $T^{*}$ vs. $\rho^{*}$ plane for the 1D case with $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$: (a) $\epsilon_{1}/\epsilon_{2}=1,0.5$, (b) $\epsilon_{1}=0$, and (c) $\epsilon_{1}/\epsilon_{2}=-0.5,-1$. Insets display the angular frequency of the asymptotic oscillations of $h(r)$ as a function of $T^{*}$ for (a) $\epsilon_{1}/\epsilon_{2}=0.5$ at $\rho^{*}=0.45$, (b) $\epsilon_{1}=0$ at $\rho^{*}=0.22$, and (c) $\epsilon_{1}/\epsilon_{2}=-1$ at $\rho^{*}=0.22$.
Since the smallest length scale of the problem is the hard core diameter $\sigma=1$, one might reasonably expect the angular frequency of the asymptotic oscillations to remain below $\omega\approx 2\pi$. However, as discussed earlier, within the loops and arcs, $\omega$ is distinctly larger than $2\pi$, indicating wavelengths significantly shorter than the hard core diameter [see the insets in Figures 1a,b and 2b,d, as well as the insets in Figure 3]. This surprising phenomenon suggests the emergence of intricate, potentially novel mesoscopic ordering that warrants deeper investigation in future studies.
2.4 $\lambda_{1}=1.35$, $\lambda_{2}=1.7$, and $\epsilon_{1}<0$: Influence of $\epsilon_{1}/\epsilon_{2}$ on FW Line
@@
We now consider the case $\epsilon_{1}<0$, where a genuine competition arises between the attractive square well with the depth $|\epsilon_{1}|$ and the repulsive barrier of the height $\epsilon_{2}$. As demonstrated in Appendix A.1, real poles of $G(s)$ may exist. If one of these real poles becomes dominant, the asymptotic decay of $h(r)$ is monotonic, and, as mentioned earlier, an FW line emerges, marking the abrupt transition between monotonic and oscillatory decay. However, a DOC line may still occur, as exemplified by Figure 3c.
@@
The results for various values of $\epsilon_{1}/\epsilon_{2}<0$ are presented in Figure 4a. A comparison of the DOC lines in Figure 3c for $\epsilon_{1}/\epsilon_{2}=-0.5$ and $-1$ with the corresponding FW lines in Figure 4a shows that the FW lines emerge at significantly lower values of $T^{*}$ and span a broader range of densities.
Figure 4: (a) FW lines on the $T^{*}$ vs. $\rho^{*}$ plane for the 1D case $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$ with, from the bottom to top, $\epsilon_{1}/\epsilon_{2}=-0.5$, $-1$, $-2$, $-4$, $-10$, and $-100$. (b) The same as in panel (a), except that now the vertical axis represents the scaled temperature $T_{1}^{*}=k_{B}T/|\epsilon_{1}|=T^{*}\epsilon_{2}/|\epsilon_{1}|$. The dotted curve is the FW line for a pure square-well fluid ($\epsilon_{2}=0$). Note that in panel (b) the curves corresponding to $\epsilon_{1}/\epsilon_{2}=-0.5$ and $-1$ are indistinguishable.
@@
The strong sensitivity of the FW lines to the values of $\epsilon_{1}/\epsilon_{2}$, as seen in Figure 4a, is significantly reduced when the temperature is scaled by the well depth $|\epsilon_{1}|$, i.e., $T_{1}^{*}=k_{B}T/|\epsilon_{1}|=T^{*}\epsilon_{2}/|\epsilon_{1}|$. This rescaling is applied in Figure 4b, which also includes the FW line for a pure square-well fluid ($\epsilon_{2}/|\epsilon_{1}|\to 0$). For the latter fluid, the FW line approaches $T_{1}^{*}\to\infty$ as $\rho^{*}\to 0$ (following a power law). However, introducing a repulsive barrier of the height $\epsilon_{2}$ causes the FW line to bend at low densities, even when $\epsilon_{2}/|\epsilon_{1}|=10^{-2}$.
It should be pointed out that in the 1D lattice model analyzed in Ref. Pȩkalski et al. (2013), the attractive interaction is limited to nearest neighbors, while the repulsion extends up to third-nearest neighbors. At $T=0$, the energy minimum is achieved by forming clusters of three consecutive particles. The authors also report the formation of clusters separated by distances greater than the range of the repulsion. However, we observe neither of these features in the exact calculations of our model. Additionally, while an FW line is identified in their work, there is no evidence of a DOC.
3 The 3D System: RFA Results
@@ 3.1 Theoretical Background
In this section, we provide a brief account of the main outcome of the RFA approach when the intermolecular potential in 3D is of the form of Equation (1). The detailed derivation may be found in References Santos et al. (2012, 2013); Yuste et al. (2022). We begin by considering a function, $G(s)$, which is distinct from its 1D counterpart. This function represents the Laplace transform of $rg(r)$; specifically,
@@
We next define an auxiliary function, $\Phi(s)$, directly related to $G(s)$ through
@@
Note that Equations (12)–(14) are formally exact. Finally, to construct our RFA, we assume the following rational function approximate form for $R_{j}(s)$:
@@
The approximation in (15) contains nine parameters to be determined by the application of certain constraints Santos et al. (2012). The expressions for those nine coefficients are presented in Appendix B.
Once again, the total correlation function $h(r)$ can be expressed in terms of the nonzero poles of $G(s)$, which, in principle, form an infinite set. These poles may be either real or occur in complex–conjugate pairs. Their locations depend on the thermodynamic state, and as before, the pole with the real part closest to zero dictates the asymptotic behavior of the total correlation function for a given state. The 3D analogs of Equations (5) and (6) are, respectively,
@@
[…]
@@
3.2 $\epsilon_{1}=\epsilon_{2}>0$: Influence of $\lambda_{2}$ on DOC Line